# A Gaussian smooth transition vector autoregressive model: An application to the macroeconomic effects of severe weather shocks

### v1

@@ 1 Introduction
In particular, we introduce a new STVAR model that we call the Gaussian smooth transition vector autoregressive (GSTVAR) model. It assumes a Gaussian conditional distribution and adopts transition weights that are similar to the mixing weights of the Gaussian mixture vector autoregressive (GMVAR) model of Kalliovirta et al. (2016). For a $p$th order model, the transition weights depend on the full distribution of the preceding $p$ observations, thereby allowing to capture complicated switching dynamics. Specifically, the greater the relative weighted likelihood of a regime is, the greater its transition weight is, which facilitates associating statistical characteristics and economic interpretations to the regimes. In contrast to the GMVAR model, which incorporates unobserved discrete regime-switches, our GSTVAR model allows for capturing gradual shifts in the dynamics of the data. The GSTVAR model also has the advantage in structural analysis that, unlike in the GMVAR model, identified structural shocks can be recovered from the fitted model.
@@
Another popular branch of the literature on nonlinear VAR models focuses on so called Markov-switching VAR (MS-VAR) models (Krolzig, 1997, see also Hamilton, 1989, 1990). They incorporate discrete regime switches that are random and unobserved, and the switching probabilities depend on the preceding regime. The MS-VAR models can flexibly capture nonlinearities in the data, but unlike our GSTVAR model, they do not accommodate gradual shifts in the dynamics nor complicated switching dynamics.
@@
We apply our GSTVAR model to study the macroeconomic effects of severe weather shocks and consider a monthly U.S. data set covering the period from 1961:1 to 2022:3 and containing an indicator of severe weather and a number of macroeconomic variables. We find a specification with two regimes sufficient. The first regime mainly dominates the earlier part of the sample period, particularly during the volatile periods of 1970’s and 1980’s, but also prevails in the later sample during the Financial crisis and the COVID-19 crisis. The second regime dominates the latter part of the sample period. Following Kim et al. (2022), we identify the severe weather shock recursively, placing the the severe weather indicator first in the vector of variables. In both regimes, a positive severe weather shock decreases GDP, consumer prices, and the interest rate, but the effects are clearly stronger in the first regime prevailing in the earlier part of the sample period and in certain crisis periods. In contrast to Kim et al. (2022), we thus conclude that the effects of the severe weather shock are stronger in the earlier than the latter part of the sample, which suggests that the U.S. economy has adapted to the changing distribution of weather related shocks over time. However, the potential adaptation may not provide sufficient resilience when the economy is in a vulnerable state.
@@
The rest of this paper is organized as follows. Section 2 first presents our framework for STVAR models, then discusses stationarity and ergodicity of the STVAR models. We also introduce the structural STVAR model and some tools useful in empirical analysis based on it. Finally, the Gaussian STVAR (GSTVAR) model is introduced by specifying the transition weights and the error distribution, and the GSTVAR model is compared to related models proposed in the previous literature. Section 3 discusses model selection and estimation of the parameters by the method of maximum likelihood. Section 4 presents the empirical application to the macroeconomic effects of severe weather shocks. Finally, Section 5 concludes. Further discussion on the stationarity condition and a Monte Carlo algorithm for estimating the generalized impulse response function (Koop et al., 1996), as well as details on the empirical application are deferred to appendices. The introduced methods are implemented to the R package sstvars (Virolainen, 2024), available on an author’s GitHub page https://github.com/saviviro/sstvars.
@@ 3.1 Maximum likelihood estimation
The true parameter value $\boldsymbol{\theta}_{0}$ is an interior point of $\boldsymbol{\Theta}$, which is a compact subset of the parameter space $\{\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{m,0},\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)\in\mathbb{R}^{M(d+d^{2}p+d(d+1)/2)}\times(0,1)^{M-1}\times(2,\infty):$ $\Omega_{m}$ is positive definite for all $m=1,...,M$, and Conditions 1 and (3.3) hold.}normal-}\}}.
@@
After the estimation rounds have been ran, the researcher can choose among the appropriate local solutions the one that maximizes the log-likelihood. Then, the sufficient Condition 1 for ergodic stationarity can be checked for the selected local solution if the more easily verified necessary Condition 2 is used to restrict the parameter space instead of Condition 1 in the estimation. The accompanying R package sstvars employs a modified genetic algorithm that works similarly to the one described in Virolainen (2022).
@@ 4.3 Impulse response analysis
To study the effects of the ACI shock, we compute the generalized impulse responses of the variables to it. As pointed out in Section 2.2, due to the endogenously determined transition weights and the fact that we allow the regime to switch as a result of a shock, there are multiple types of possible asymmetries. Therefore, we compute generalized impulse responses that can vary depending on the initial values of the variables as well as on the sign and size of the shock. We are particularly interested in finding out about the potential state-dependence of the effects of the ACI shock, and to capture them, we draw initial values from the stationary distribution of each regime separately. Shocks of different size seem to result in similar impulse responses, so we only report results for one standard-error shocks. The GIRFs are scaled so that the instantaneous response of the ACI index is $0.3$, making the responses to shocks of different signs and sizes comparable.
Figure 2: Generalized impulse response functions to the identified ACI shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of the ACI index, GDP growth rate, inflation growth rate, the interest rate variable, and transition weights of Regime 1 are depicted in each row. The left column shows the responses to a positive (grey solid line) and negative (black dashed line) one-standard-error shocks with initial values generated from the stationary distribution of Regime 1. The right column shows the corresponding responses in Regime 2. All GIRFs have been scaled so that the instantaneous change in the ACI index is $0.3$. For a positive ACI shock, the $95\%$ confidence intervals are depicted with shaded grey areas, whereas for a negative ACI shock, they are depicted as black dashed lines. The results are based on $2500$ draws of initial values (for details of the algorithm, see Appendix B).
Figure 2 presents the GIRFs (and the 95% confidence intervals) of the variables, including the transition weights, to the identified ACI shock $h=0,1,...,36$ quarters ahead. The right and left columns depict the responses to a one-standard-error positive and negative shocks in Regime 1 and Regime 2, respectively. In both regimes, a positive ACI shock mainly decreases (and a negative shock increases) GDP, consumer prices, and the interest rate. However, there is a major difference in the GIRFs between the regimes: the effects of the ACI shock on all macroeconomic variables are clearly stronger in Regime 1. This may be the case because this regime appears to be attributed to more volatile times, including recessions and crisis periods. Thus, it might represent a state of the economy that is more vulnerable to weather related shocks, and therefore, a similar shock is likely to result in greater damage in Regime 1 than in Regime 2. Other differences between the regimes are small.
@@ 5 Conclusion
By making use of the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we presented a sufficient condition for stationarity and ergodicity of the GSTVAR model. We also discussed estimation of the model parameters by the method of maximum likelihood, which, due the complexity of the log-likelihood function, can be very tedious in practice. Hence, following Meitz et al. (2023) and Virolainen (2022, forthcoming), we proposed employing a two-phase procedure in which a genetic algorithm is used to find starting values for a gradient based method. The introduced methods, including a modified version of a genetic algorithm, are implemented to the R package sstvars, available on an author’s GitHub page https://github.com/saviviro/sstvars.
@@ C.1 Model selection and adequacy of the selected model
Figure C.1: Partial autocorrelation functions for the first $24$ lags estimated for the monthly U.S. series covering the period from 1961I to 2022III.
@@
Figure C.2: Auto- and crosscorrelation functions of the residuals of the fitted two-regime GSTVAR $p=4$ model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The blue dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ ($T=731$ as the first $p=4$ observations were used as the initial values) for autocorrelations of IID observations.
Figure C.3: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime fourth-order GSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The blue dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ ($T=731$ as the first $p=4$ observations were used as the initial values) for autocorrelations of IID observations.
@@ C.2 Robustness checks
Figure C.5: Estimated transition weights of the first regime for the two-regime fourth-order GSTVAR model fitted to the monthly U.S. data covering the period from 1963:4 to 2019:5. The shaded areas indicate the period of NBER based U.S. recessions.
@@
The GIRFs estimated for the model are presented in Figure C.6, which is similar to Figure 2 of the main paper. The GIRFs are mostly similar to our main specification utilizing the full sample period from 1961:1 to 2022:3. In both of the regimes, a positive ACI shock decreases GDP and consumer prices, but the effects are stronger in Regime 1. The GIRFs are close to linear in Regime 1, and while there is variation in the GIRFs depending on the initial values in Regime 2, the effects of positive and negative shocks (as well as the effects small and large shocks not shown in Figure C.6) are very similar. A notable difference to our main specification is that a positive ACI shock increases the interest rate in Regime 1 but decreases it in Regime 2.
Figure C.6: Generalized impulse response functions to the identified ACI shock $h=0,1,....,36$ months ahead, based on the data from from 1963:4 to 2019:5. From the top to bottom, the responses of the ACI index, GDP growth rate, inflation growth rate, the interest rate variable, and transition weights of Regime 1 are depicted in each row, respectively. The left column shows the responses to positive (grey solid line) and negative (back dashed line) one-standard-error shocks with initial values generated from the stationary distribution of Regime 1. All GIRFs have been scaled so that instantaneous effect of the ACI index is $0.3$. For a positive ACI shock, the $95\%$ confidence intervals that reflect uncertainty about the initial state within the given regime are depicted with shaded grey areas, whereas for a negative ACI shock, they are depicted with black dashed lines. The results are based on 2500 draws of initial values (see Appendix B for details).

### v2

@@ 1 Introduction
In particular, we introduce a new STVAR model that we call the Gaussian smooth transition vector autoregressive (GSTVAR) model. It assumes a Gaussian conditional distribution and adopts transition weights that are similar to the mixing weights of the Gaussian mixture vector autoregressive (GMVAR) model of Kalliovirta et al. (2016). For a $p$th order model, the transition weights depend on the full distribution of the preceding $p$ observations, thereby allowing to capture complicated switching dynamics. Specifically, the greater the relative weighted likelihood of a regime is, the greater its transition weight is, which facilitates associating statistical characteristics and economic interpretations to the regimes. In contrast to the GMVAR model, which incorporates unobserved discrete regime switches, our GSTVAR model allows for capturing gradual shifts in the dynamics of the data. The GSTVAR model also has the advantage in structural analysis that, unlike in the GMVAR model, identified structural shocks can be recovered from the fitted model.
@@
Another popular branch of the literature on nonlinear VAR models focuses on so-called Markov-switching VAR (MS-VAR) models (Krolzig, 1997, see also Hamilton, 1989, 1990). They incorporate discrete regime switches that are random and unobserved, and the switching probabilities depend on the preceding regime. The MS-VAR models can flexibly capture nonlinearities in the data, but unlike our GSTVAR model, they do not accommodate gradual shifts in the dynamics nor complicated switching dynamics.
In addition to nonlinear VAR models, time-varying parameter VAR (TVP-VAR) models (Cogley and Sargent, 2001; Cogley and Sargent, 2005; Primiceri, 2005; Koop et al., 2009; among others) allow for capturing random but persistent time-variation in parameter values. However, compared to our GSTVAR model, they have the major limitation that the stochastic processes governing the changes in parameter values are exogenous to the variables included. Consequently, changes in the macroeconomic variables have no influence on how the parameter values evolve, and, thus, potentially important endogenous dynamics of the economy are overlooked.
@@
We apply our GSTVAR model to study the macroeconomic effects of severe weather shocks and consider a monthly U.S. data set covering the period from 1961:1 to 2022:3 and containing an indicator of severe weather and a number of macroeconomic variables. We find a specification with two regimes sufficient. The first regime mainly dominates the earlier part of the sample period, particularly during the volatile periods of 1970’s and 1980’s, but also prevails in the later sample during the Financial crisis and the COVID-19 crisis. The second regime dominates the latter part of the sample period. Following Kim et al. (2022), we identify the severe weather shock recursively, placing the severe weather indicator first in the vector of variables. In both regimes, a positive severe weather shock decreases GDP, consumer prices, and the interest rate, but the effects are clearly stronger in the first regime prevailing in the earlier part of the sample period and in certain crisis periods. In contrast to Kim et al. (2022), we thus conclude that the effects of the severe weather shock are stronger in the earlier than the latter part of the sample, which suggests that the U.S. economy has adapted to the changing distribution of weather related shocks over time. However, the potential adaptation may not provide sufficient resilience when the economy is in a vulnerable state.
@@
The rest of this paper is organized as follows. Section 2 first presents our framework for STVAR models, then discusses stationarity and ergodicity of the STVAR models. We also introduce the structural STVAR model and some tools useful in empirical analysis based on it. Finally, the Gaussian STVAR (GSTVAR) model is introduced by specifying the transition weights and the error distribution, and the GSTVAR model is compared to related models proposed in the previous literature. Section 3 discusses model selection and estimation of the parameters by the method of maximum likelihood. Section 4 presents the empirical application to the macroeconomic effects of severe weather shocks. Finally, Section 5 concludes. Further discussion on the stationarity condition and a Monte Carlo algorithm for estimating the generalized impulse response function (Koop et al., 1996), as well as details on the empirical application are deferred to appendices. The introduced methods are implemented to the accompanying R package sstvars (Virolainen, 2024), available in the CRAN repository.
@@ 2.3.1 Comparison to related models
Besides nonlinear VAR models, time-varying parameter VAR (TVP-VAR) models (Cogley and Sargent, 2001; Cogley and Sargent, 2005; Primiceri, 2005; Koop et al., 2009; among others) offer a flexible framework for capturing time-variation in the parameter values that is quite different from our GSTVAR model. The standard deviations of the reduced form innovations are often assumed to evolve as geometric random walks (i.e., they accommodate stochastic volatility), and the rest of the parameter values are typically either constant or evolve as random walks (possibly with mixture innovations, see, e.g., Koop et al., 2009). TVP-VAR models are very capable of capturing gradual as well as abrupt changes in parameter values, but they have the major limitation that the stochastic processes governing the parameters are exogenous to the endogenous variables. Hence, although the parameter estimates can reflect changes in economic conditions, the included variables do not directly influence the evolution of the parameter values, and, thus, potentially important endogenous dynamics of the economy are ignored. In contrast, our GSTVAR model allows the included variables to affect the transition weights at each point of time through the full distribution of the preceding $p$ observations.
@@ 3.1 Maximum likelihood estimation
The true parameter value $\boldsymbol{\theta}_{0}$ is an interior point of $\boldsymbol{\Theta}$, which is a compact subset of the parameter space $\{\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{m,0},\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)\in\mathbb{R}^{M(d+d^{2}p+d(d+1)/2)}\times(0,1)^{M-1}\times(2,\infty):$ $\Omega_{m}$ is positive definite for all $m=1,...,M$, and Conditions 1 and (3.3) hold.}}\}}.
@@
After the estimation rounds have been ran, the researcher can choose among the appropriate local solutions the one that maximizes the log-likelihood. Then, the sufficient Condition 1 for ergodic stationarity can be checked for the selected local solution if the more easily verified necessary Condition 2 is used to restrict the parameter space instead of Condition 1 in the estimation. The accompanying R package sstvars (Virolainen, 2024) employs a modified genetic algorithm that works similarly to the one described in Virolainen (2022).
@@ 4.3 Impulse response analysis
To study the effects of the ACI shock, we compute the generalized impulse responses of the variables to it. As pointed out in Section 2.2, due to the endogenously determined transition weights and the fact that we allow the regime to switch as a result of a shock, there are multiple types of possible asymmetries. Therefore, we compute generalized impulse responses that can vary depending on the initial values of the variables as well as on the sign and size of the shock. We are particularly interested in finding out about the potential state-dependence of the effects of the ACI shock, and to capture them, we draw initial values from the stationary distribution of each regime separately. Shocks of different sign and size seem to result in similar impulse responses, so we only report results for a one standard-error positive shock. The GIRFs are scaled so that the instantaneous response of the ACI index is $0.3$, making the responses comparable.
Figure 2: Generalized impulse response functions to the identified ACI shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of the ACI index, GDP growth rate, inflation growth rate, the interest rate variable, and transition weights of Regime 1 are depicted in the rows. The left column shows the responses to a positive one-standard-error shock with initial values generated from the stationary distribution of Regime 1. The right column shows the corresponding responses in Regime 2. All GIRFs have been scaled so that the instantaneous change in the ACI index is $0.3$. The $95\%$ confidence intervals are depicted with shaded grey areas. The results are based on $2500$ draws of initial values (for details of the algorithm, see Appendix B).
Figure 2 presents the GIRFs (and the 95% confidence intervals) of the variables, including the transition weights, to the identified ACI shock $h=0,1,...,36$ quarters ahead. The right and left columns depict the responses to a one-standard-error positive shock in Regime 1 and Regime 2, respectively. In both regimes, a positive ACI shock mainly decreases GDP, consumer prices, and the interest rate. However, there is a major difference in the GIRFs between the regimes: the effects of the ACI shock on all macroeconomic variables are clearly stronger in Regime 1. This may be the case because this regime appears to be attributed to more volatile times, including recessions and crisis periods. Thus, it might represent a state of the economy that is more vulnerable to weather related shocks, and therefore, a similar shock is likely to result in greater damage in Regime 1 than in Regime 2. Other differences between the regimes are small.
@@ 5 Conclusion
By making use of the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we have presented a sufficient condition for stationarity and ergodicity of the GSTVAR model. We have also discussed estimation of the model parameters by the method of maximum likelihood, which, due the complexity of the log-likelihood function, can be very tedious in practice. Hence, following Meitz et al. (2023) and Virolainen (2022, forthcoming), we propose employing a two-phase procedure in which a genetic algorithm is used to find starting values for a gradient based method. The introduced methods, including a modified version of a genetic algorithm, are implemented in the CRAN distributed R package sstvars (Virolainen, 2024). A potential area of future research is an extension of our model that accommodates strong forms of conditional heteroskedasticity by incorporating autoregressive conditional heteroskesticity or stochastic volatility to the shocks.
@@ C.1 Model selection and adequacy of the selected model
Figure C.1: Partial autocorrelation functions for the first $24$ lags estimated for the monthly U.S. series covering the period from 1961:1 to 2022:3.
@@
Finally, we assess whether the regime shifts can be attributed solely to changes in the level or volatility of the series, or whether the data exhibit time-variation also in the autoregressive dynamics. Since our GSTVAR model is ergodic stationary, the results of consistency and conventional limiting distribution apply to the ML estimator under conventional high level conditions, such that standard likelihood based tests are available. Hence, to check if the regime shifts are driven only by changes in the volatility regime, we test whether the intercepts and AR matrices are identical across the regimes (i.e., time-invariant). The Wald test for the null of constancy of intercepts and AR matrices produces a $p$-value smaller than $10^{-7}$, indicating that the regime-shifts are not solely driven by changes in the volatility regime. The Wald test for the constancy of the AR matrices alone produces a $p$-value smaller than $10^{-5}$. Therefore, we conclude in favor of regime shifts in the level, volatility, as well as the autoregressive dynamics of the series.
Figure C.2: Auto- and crosscorrelation functions of the residuals of the fitted two-regime GSTVAR $p=4$ model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
Figure C.3: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime fourth-order GSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@ C.2 Robustness checks
Figure C.5: Estimated transition weights of Regime 1 for the two-regime fourth-order GSTVAR model fitted to the monthly U.S. data covering the period from 1963:4 to 2019:5. The shaded areas indicate the period of NBER based U.S. recessions.
@@
The GIRFs estimated for the model are presented in Figure C.6, which is similar to Figure 2 of the main paper. Responses to shocks of different sign and size seem similar, so only the responses to a positive one-standard-error shock are reported. The GIRFs are mostly similar to our main specification utilizing the full sample period from 1961:1 to 2022:3. In both of the regimes, a positive ACI shock decreases GDP and consumer prices, but the effects are stronger in Regime 1. The GIRFs are close to linear in Regime 1, while there is variation in the GIRFs depending on the initial values in Regime 2. A notable difference to our main specification is that a positive ACI shock increases the interest rate in Regime 1 but decreases it in Regime 2.
Figure C.6: Generalized impulse response functions to the identified ACI shock $h=0,1,....,36$ months ahead, based on the data from from 1963:4 to 2019:5. From the top to bottom, the responses of the ACI index, GDP growth rate, inflation growth rate, the interest rate variable, and transition weights of Regime 1 are depicted in the rows. The left column shows the responses to a positive one-standard-error shocks with initial values generated from the stationary distribution of Regime 1. The right column shows the corresponding responses in Regime 2. All GIRFs have been scaled so that instantaneous effect of the ACI index is $0.3$. The $95\%$ confidence intervals that reflect uncertainty about the initial state within the given regime are depicted with shaded grey areas. The results are based on 2500 draws of initial values (see Appendix B for details).